# Mantis: Interleaved Multi-Image Instruction Tuning

### v2

@@ 3.3 Results on Multi-Image Tasks
Model Size TextVQA VQA MMB MMMU OKVQA SQA MathVista Avg
OpenFlamingo 9B 46.3 58.0 32.4 28.7 51.4 45.7 18.6 40.2
Idefics1 9B 39.3 68.8 45.3 32.5 50.4 51.6 21.1 44.1
InstructBLIP 7B 33.6 75.2 38.3 30.6 45.2 70.6 24.4 45.4
Yi-VL 6B 44.8 72.5 68.4 39.1 51.3 71.7 29.7 53.9
Qwen-VL-Chat 7B 63.8 78.2 61.8 35.9 56.6 68.2 15.5 54.3
LLaVA-1.5 7B 58.2 76.6 64.8 35.3 53.4 70.4 25.6 54.9
@@
Mantis-CLIP 8B 56.4 73.0 66.0 38.1 53.0 73.8 31.7 56.0
Mantis-SigLIP 8B 59.2 74.9 68.7 40.1 55.4 74.9 34.4 58.2
@@
We report our results on the multi-image task in  Table 4. We put the GPT-4V’s results at the top of the table as it’s a closed-source LMM. We can see that our best version Mantis-Idefics2 already matches the overall performance of GPT-4V. We put comparison results between Mantis and other open-source LMMs, along with the insights derived in the following paragraphs.
Mantis performs well on held-in evaluation We found that Mantis-SigLIP without any multi-image pre-training can already achieve best known performance on these two benchmarks. Mantis-Idefics2 has better performance of $89.71$ and $75.20$ on NLVR2 and Q-Bench respectively, marking it the SoTA for these two tasks. Notably, Mantis-Idefics2 also beats the Idefics2 by $2.84$ points, which was already trained on NLVR2. It shows the cross-task benefits of our instruction tuning. The remarkable performance on these 2 benchmarks demonstrate that Mantis has gained strong ability in multi-image logical reasoning and low-level vision understanding.
Mantis generalizes well on held-out evaluation We show that our models can attain strong performance on other held-out benchmarks, Mantis-Eval, BLINK (val), and MVBench. Mantis-SigLIP achieves $59.45$ on Mantis-Eval, while Mantis-Idefics2 achieves $49.05$, and $51.38$ on BLINK and MVBench respectively, surpassing all other baselines on these held-out tasks. The gain over the existing strongest baseline is particularly tremendous on Mantis-Eval and MVBench.
Multi-Image Pre-training is not necessary In these experiments, Mantis-CLIP and Mantis-SigLIP are attaining much better performance than Mantis-Flamingo and similar performance as Mantis-Idefics. These results reveal that the multi-image pre-training (on massive web corpus) is not a necessary path toward strong multi-image performance.
‘Merge’ vs ‘Sequence’ input The ‘merge’ evaluation is an important baseline to evaluate the multi-image understanding ability of those LMMs that can only accept one image at a time. LLaVA-v1.6’s dynamic image resolution feature can divide a single image into multiple tiles, processing the ‘merge’ image like multiple images [7]. But it still performs poorly on multi-image tasks, only getting $45.62$ and $39.55$ on the Mantis-Eval and BLINK benchmark, which are at least 10 absolute points lower than Mantis. This result indicates the advantages of building models to accept sequence image inputs.
@@ 3.4 Results on Single-Image Tasks
We also evaluate Mantis-CLIP and Mantis-SigLIP on various single-image tasks, including TextVQA [41], VQA-v2 [42], MMBench [43], MMMU [44], etc. Results are shown in  Table 5. All the evaluations are conducted with the help of LMMs-Eval [45] tool.
Compared to a set of popular LMMs, Mantis-SigLIP gets significant improvements on MMBench-English, MMMU, and ScienceQA benchmarks though we did not specifically optimize the single-image abilities. Solving problems in these tasks requires not only the visual perception ability but also the good reasoning ability of the LLM backbone. Considering that we are using the powerful LLaMA-3 as the backbone LLM, these improvements are expected. We also compare the results of Mantis-Idefics2 against Idefics2 on single-image tasks. On several tasks like MMB, MMMU, and OKVQA, Mantis-Idefics maintains a similarly strong performance as Idefics2. However, we do observe some drops in ScienceQA, MathVista, which are mainly due to a lower mix ratio of OCR and Math data in our instruction tuning. Overall, the average drop of 4% is in the tolerable range.
@@ 3.5 Ablation Studies
Is multi-image instruction tuning necessary? To investigate, we further train a Mantis-Flamingo based on OpenFlamingo [14], which has been trained on MMC4 [46], a text-image interleaved pre-training dataset. As shown in Table 4, the average performance of Mantis-Flamingo increased by $19.2$ points after further fine-tuning on Mantis-Instruct. The results show that there is still a huge improvement space, even if a model is fully pre-trained on a multi-image dataset, which demonstrates the necessity of learning multi-image skills in the fine-tuning phase.
@@
Table 6: Ablation on whether multi-image pre-training is necessary. One model is pre-trained on the CC3M subset, while the other is additionally pre-trained on the a OBELICS-100K multi-image dataset. They are both fine-tuned on a subset Mantis-Instruct for comparison. Results show minor improvement with additional multi-image pre-training.
@@
Table 7: Data ablation study results of different subsets of Mantis-Instruct, where each skill corresponding to a specific subsets defined in Table 1. $\Delta$ column of each row represents the performance improvements after adding new subsets to the last column results.
@@ 4.1 Large Multimodal Models
Large Multimodal Models (LMMs) are mostly denoted as large language models that can understand multiple modalities besides the human language. Some works aim to fuse image, audio, video, or more modalities with language [47, 48], while more works mainly focus on better-combining vision knowledge with the language [49, 14, 36]. BLIP-2 has curated a large-scale image captioning dataset, bootstrapping a language model along with a vision encoder to be a powerful multimodal model [8]. Following this, LLaVA further proposes a cheaper approach to train a powerful LMM through visual instruction tuning [32]. LLaVA-Next further optimizes single-image performance, with a cost of an increasing number of image tokens per token [7], consuming more than 2k for each image, which is about 4 times of the original LLaVA model. Later LMMs like QwenVL [31], CogVLM [9], Yi-VL [40], etc. all follow a similar architecture of LLaVA. However, real-world visual problems are way more complex and usually require a sequence of images to describe the situation. Although better performance is achieved in the single-image scenario, it’s doubtful whether it’s worth the cost.
@@ 4.2 Multi-Image Supported Large Multimodal Models
Previous works have also noticed the importance of multi-image ability. Deepmind’s close-sourced Flamingo [49] focuses on optimizing in-context learning for a sequence of image/text pairs but lacking free-from image-text interleaved training examples. Kosmos2 [33], Fuyu [25] both support text-image interleaved format given the model structure, but they did not optimize in multi-image reasoning, and the released inference codes also do not support multi-image as input. Emu2 [15] is a generative multimodal model that supports interleaved text-image inputs, as well as generating both images and texts. However, it’s mainly focused on in-context examples like Flamingo and auto-regressive image generation, instead of the 4 skills mentioned in  section 1 Another common variant of LMMs that can accept multiple images as input is video understanding models, such as Video-LLaMA [50]. However, as shown in the MVBench benchmark, it is also shown that Video-LLaMA behaves worse than LMMs that can only accept one image as input [30], which raises doubts about its multi-image understanding ability. Different from these works, Mantis aims to optimize the multi-image ability of LMMs guided by the 4 defined skills, co-reference, reasoning, comparison, and temporal understanding, thus equipping the model with better multi-image ability.
@@ 4.3 Multi-Image Evaluation
While many benchmarks have been developed for single-image tasks, such as TextVQA [41], VQA-V2 [42], MMBench [43], and so on, the multi-image evaluation did not capture much attention until recently. Difference description is the first task which large-scale data and good evaluation benchmarks are developed for, including works like Spot-the-Diff [19], Birds-to-Words [21], etc. They are usually evaluated by n-gram metrics like BLEU, and ROUGE, and only involve two images. The later NLVR2 [18] further includes logical reasoning across 2 images into evaluation, focusing on a higher level of cognition ability. Q-bench [28] and BLINK [29] are both proposed for low-level multi-image vision tasks, where humans easily do but model struggles. Another set of multi-image evaluations comes from video or image sequence evaluation, including Mementos [51], MVBench [30], etc. In this case, the model usually needs to take more than 2 images to answer. Recently, there are also attempts to prompt LMMs to evaluate vision-language task [52], or image-generation tasks [53], which also demands good multi-image reasoning ability. The accuracies of the these tasks of different LMMs can be seen as another form of benchmark for multi-image ability.
@@ A.1 Details of Mantis-Instruct subsets
Co-Instruct [35]: Co-Instruct is a dataset curated to compare the quality of 2-4 images in the format of open-ended answer or multi-choice QA format. It’s similar to the tasks in the Qbench [28].
@@
NLVR2 [18]: NLVR2 is a natural language reasoning dataset grounded in images. The task is to determine whether a statement is true or false given the visual contents of 2 images. NLVR2 covers many linguistic phenomena, such as cardinality, existential quantifiers, universal quantifiers, etc.
@@
Dreamsim [20]: DreamSim proposes a dataset to train a model to judge image similarity. We reuse their datasets, and the task is to judge which candidate image is more similar to a reference image.
@@
ImageCoDe [54]: ImageCoDe is a multimodal challenge called image retrieval from contextual descriptions, where a multimodal model is required to retrieve/select the correct image from a set of 10 minimally contrastive images. For each image, 9 similar images are retrieved from Open Images Dataset V6 [55] using CLIP encodings.
@@
Contrast-Caption: Inspired by the paradigm of contrastive learning, we reformat existing image captioning tasks into a multi-image version, where 2 tasks are defined. Given multiple images, the first task we defined is to judge which image matches the provided caption, and the second task is to generate a caption for a denoted image. We randomly select 2 to 8 images along with their captions to form a single item, then apply the template of the 2 tasks defined. We used the images and captions provided by ShareGPT-4V [56] and LAION GPT-4V captions from LVIS [57].
@@
Spot-the-diff [19]: This is a dataset that requires the model to generate the difference description given 2 images. We use the processed version from MIMIC-IT [58].
@@
LRV-multi [59]: This is a single-image dataset used to mitigate hallucinations of LMMs. We process it similarly to LLaVA-665k-multi to get a "synthetic" multi-image dataset.
@@
Birds-to-Words [21]: The task of this dataset is to generate different descriptions and give 2 images of different birds. We directly take the training split for training. Besides, we prompt GPT-3.5-turbo to generate a possible question given the reference answer, so we can get a diverse question pool.
Multi-VQA: This dataset contains multi-image reasoning QA pairs, where each question is generated by GPT-4 based on the captions of the images. Each question is required to involve at least two images. Images and captions are sampled from ShareGPT-4V [56]. The prompt template is in  Table 8.
@@
Video QA: We also include some video question answering datasets, including NExTQA [22], STAR [23], and visual-story-telling [32]. nextqa and star are the processed version from Flipped-VQA [60]. The visual-story-telling are the processed version from MIMIC-IT [58].
@@
Single-image-VQA: Instead of focusing only on multi-image tasks, we claim that it’s necessary to also include some single-image VQA dataset to avoid catastrophic forgetting on single-image tasks. We included some of the single image data from LLaVA-665k as well as DocVQA [61], DVQA [62], and ChartQA [63] to enhance its ability on diagrams and OCR. In practice, since the size of DVQA is too large, we only use 30k out of 200k examples for the training.
A.2 Heuristics of data curation
Similar to visual instruction tuning [5], data curation is an important step to ensure good performance. During the curation, we have found some heuristics that can improve the performance. We list them below.
  1. 1.
    Conflicts between short answers and long answers. Many datasets we collected are domain-specific, and some of them are classification tasks like NLVR2 [18] and DreamSim [20]. For these datasets, we convert them into multiple-choice QA format to avoid conflicts between repeated short-answers and some long answers from other datasets.
  2. 2.
  3. 3.
    Positions of the "<image>" placeholder. For each item with multiple images, we write some rules to randomly put the image placeholders at the beginning of the first question or the end of the first question. This technique is also used by the curation LLaVA-665k [6].
A.3 Training Details
For Mantis-CLIP and Mantis-SigLIP, we first pre-train the multimodal projector on LLaVA pre-train data, where the learning rate is set to 1e-3, the batch size is set to 256. After pre-training, we fine-tune the 2 models on the Mantis-Instruct. For Mantis-Flamingo and Mantis-Idefics2, we directly fine-tune them on Mantis-Instruct since they have already been pre-trained on millions of data.
During the fine-tuning, we train each model on the data for 1 epoch, with a batch size of 128. The maximum context length is set to 8192. The learning rate is all set to 1e-5, except for Idefics2, where the learning rate is set to 5e-6 to better preserve its original knowledge. We set the warmup ratio to be 0.03 and use a cosine learning rate scheduler. We speed up our training and inference with Flash-attention2 [64]. We use DeepSpeed Zero-3 [65] for full fine-tuning. We apply QLoRA [66] along with DoRA [67] to more efficiently do comprehensive ablation studies under limited resources. All full fine-tuning ran on 16 A100 GPUs while the ablation study ran on 8 A100 GPUs.
A.4 Prompt Template
Table 8: Propmt template used to curate the Multi-VQA subset.
Output format:
Question: First question?
Answer: The answer to the first question.
Question: Second question?
Answer: The answer to the second question.

### v3

@@ 3.4 Results on Multi-Image Tasks
Model Size TextVQA VQA MMB MMMU OKVQA SQA Avg
OpenFlamingo 9B 46.3 58.0 32.4 28.7 51.4 45.7 43.8
Idefics1 9B 39.3 68.8 45.3 32.5 50.4 51.6 48.0
InstructBLIP 7B 33.6 75.2 38.3 30.6 45.2 70.6 48.9
Yi-VL 6B 44.8 72.5 68.4 39.1 51.3 71.7 58.0
Qwen-VL-Chat 7B 63.8 78.2 61.8 35.9 56.6 68.2 60.8
LLaVA-1.5-Vicuna 7B 58.2 76.6 64.8 35.3 53.4 70.4 59.8
LLaVA-1.6-Vicuna 7B 64.9 81.8 67.4 35.8 44.0 68.5 60.4
@@
Mantis-CLIP 8B 56.4 73.0 66.0 38.1 53.0 73.8 60.1
Mantis-SigLIP 8B 59.2 74.9 68.7 40.1 55.4 74.9 62.2
@@
We report our results on the multi-image task in  Table 5. We put the results of GPT-4V/o at the top of the table as it’s a closed-source LMM. We can see that our best version Mantis-Idefics2 almost matches the overall performance of GPT-4V. We put comparison results between Mantis and other open-source LMMs, along with the insights derived in the following paragraphs.
Mantis performs well on held-in evaluation: We found that Mantis-SigLIP without any multi-image pre-training can already achieve the best-known performance on these two benchmarks. Mantis-Idefics2 has better performance of $89.71$ and $75.20$ on NLVR2 and Q-Bench respectively, marking it the SoTA for these two tasks. Notably, Mantis-Idefics2 also beats the Idefics2 by $2.84$ points, which was already trained on NLVR2. It shows the cross-task benefits of our instruction tuning. The remarkable performance on these 2 benchmarks demonstrates that Mantis has gained strong ability in multi-image logical reasoning and low-level vision understanding.
Mantis generalizes well on held-out evaluation: We show that our models can attain strong performance on the 6 held-out benchmarks. Mantis-SigLIP achieves $59.45$ on Mantis-Eval, and Mantis-Idefics2 achieves $49.05$, and $51.38$ on BLINK and MVBench respectively, surpassing all other baselines on these held-out tasks. It’s also worth mentioning the reported performance of Mantis on the 3 recently released multi-image benchmarks with comprehensive evaluation tasks. Mantis models surpassed the VILA and Idefics2, which are pre-trained with millions of multi-image data, by a large margin, making them the best open-source models on the leaderboard. This is strong evidence of Mantis’s good generalization ability in the unseen scenario.
Multi-Image Pre-training is not necessary: In these experiments, Mantis-CLIP and Mantis-SigLIP are attaining much better performance than Mantis-Flamingo, and similar performance as Mantis-Idefics, which are pre-trained on millions of text-image interleaved data. These results reveal that the multi-image pre-training (on massive web corpus) is not a necessary path toward strong multi-image performance.
‘Merge’ vs ‘Sequence’ input: The ‘merge’ evaluation is an important baseline to evaluate the multi-image understanding ability of those LMMs that can only accept one image at a time. LLaVA-v1.6’s dynamic image resolution feature can divide a single image into multiple tiles, processing the ‘merge’ image like multiple images (Liu et al., 2024). But it still performs poorly on multi-image tasks, only getting $45.62$ and $39.55$ on the Mantis-Eval and BLINK benchmark, which are at least 10 absolute points lower than Mantis. This result indicates the advantages of building models to accept sequence image inputs.
Vision encoder matters: Unlike many other LMMs like Flamingo, LLaVA, and BLIP, Fuyu is a decoder-only LMM without a vision encoder. To investigate the importance of vision encoder, we also train an additional Mantis-Fuyu as a comparison. Results in  Table 5 show that the performance of Mantis-Fuyu is worse than Mantis-CLIP and Mantis-SigLIP where a vision encoder exists. While we acknowledge a possible factor brought by the different LLM backbones, we think whether vision encoder exists is a bigger difference and conclude that a good vision encoder can enhances the performance.
@@ 3.5 Results on Single-Image Tasks
We also evaluate Mantis-CLIP and Mantis-SigLIP on various single-image tasks, including TextVQA (Singh et al., 2019a), VQA-v2 (Goyal et al., 2016), MMBench (Liu et al., 2023d), MMMU (Yue et al., 2023), etc. Results are shown in  Table 6. All the evaluations are conducted with the help of LMMs-Eval (Li* et al., 2024) tool. Mantis-Flamingo’s and Mantis-Fuyu’s performance is omitted due to its bad performance.
Compared to a set of popular LMMs, Mantis-SigLIP gets significant improvements on MMBench-English, MMMU, and ScienceQA benchmarks though we did not specifically optimize the single-image abilities. Solving problems in these tasks requires not only the visual perception ability but also the good reasoning ability of the LLM backbone. Considering that we are using the powerful LLaMA-3 as the backbone LLM, these improvements are expected. We also compare the results of Mantis-Idefics2 against Idefics2 on single-image tasks. On several tasks like MMB, MMMU, and OKVQA, Mantis-Idefics maintains a similarly strong performance as Idefics2. However, we do observe some drops in ScienceQA, which are mainly due to a lower mix ratio of OCR and Math data in our instruction tuning. Overall, the average drop of 4% is tolerable.
@@ 3.6 Ablation Studies
Is multi-image instruction tuning necessary? To investigate, we further train a Mantis-Flamingo based on OpenFlamingo (Awadalla et al., 2023), which has been trained on MMC4 (Zhu et al., 2023), a text-image interleaved pre-training dataset. As shown in Table 5, the average performance of Mantis-Flamingo increased by $19.2$ points after further fine-tuning on Mantis-Instruct. The results show that there is still a huge improvement space, even if a model is fully pre-trained on a multi-image dataset, which demonstrates the necessity of learning multi-image skills in the fine-tuning phase.
@@
Table 7: Ablation on whether multi-image pre-training is necessary. One model is pre-trained on the CC3M subset, while the other is additionally pre-trained on the a OBELICS-100K multi-image dataset. They are both fine-tuned on a subset Mantis-Instruct for comparison. Results show minor improvement with additional multi-image pre-training. Each ablation model is trained with QLoRA with DoRA.
@@
Table 8: Data ablation study results of different subsets of Mantis-Instruct, where each skill corresponding to a specific subsets defined in Table 1. $\Delta$ column of each row represents the performance improvements after adding new subsets to the last column results. Each ablation model is trained with QLoRA with DoRA.
@@ 4.1 Large Multimodal Models
Large Multimodal Models (LMMs) are mostly denoted as large language models that can understand multiple modalities besides the human language. Some works aim to fuse image, audio, video, or more modalities with language (Wu et al., 2023b; Zhan et al., 2024), while more works mainly focus on better-combining vision knowledge with the language (Alayrac et al., 2022; Awadalla et al., 2023; Dai et al., 2023). BLIP-2 has curated a large-scale image captioning dataset, bootstrapping a language model along with a vision encoder to be a powerful multimodal model (Li et al., 2023c). Following this, LLaVA further proposes a cheaper approach to train a powerful LMM through visual instruction tuning (Huang et al., 2016). LLaVA-Next further optimizes single-image performance, with a cost of an increasing number of image tokens per token (Liu et al., 2024), consuming more than 2k for each image, which is about 4 times of the original LLaVA model. Later LMMs like QwenVL (Bai et al., 2023), CogVLM (Wang et al., 2023), Yi-VL (AI et al., 2024), etc. all follow a similar architecture of LLaVA. However, real-world visual problems are way more complex and usually require a sequence of images to describe the situation. Although better performance is achieved in the single-image scenario, it’s doubtful whether it’s worth the cost.
@@ 4.2 Multi-Image Supported Large Multimodal Models
Previous works have also noticed the importance of multi-image ability. Deepmind’s close-sourced Flamingo (Alayrac et al., 2022) focuses on optimizing in-context learning for a sequence of image/text pairs but lacking free-from image-text interleaved training examples. Kosmos2 (Peng et al., 2023), Fuyu (Adept AI, 2023) both support text-image interleaved format given the model structure, but they did not optimize in multi-image reasoning, and the released inference codes also do not support multi-image as input. Emu2 (Sun et al., 2023) is a generative multimodal model that supports interleaved text-image inputs, as well as generating both images and texts. However, it’s mainly focused on in-context examples like Flamingo and auto-regressive image generation, instead of the 4 skills mentioned in  section 1 Another common variant of LMMs that can accept multiple images as input is video understanding models, such as Video-LLaMA (Zhang et al., 2023a). However, as shown in the MVBench benchmark, it is also shown that Video-LLaMA behaves worse than LMMs that can only accept one image as input (Li et al., 2023d), which raises doubts about its multi-image understanding ability. Different from these works, Mantis aims to optimize the multi-image ability of LMMs guided by the 4 defined skills, co-reference, reasoning, comparison, and temporal understanding, thus equipping the model with better multi-image ability. Our contemporary works on optimizing interleaved multi-image ability include LLaVA-Next-Interleave (Li et al., 2024b) and LLaVA-OneVision (Li et al., 2024a), which also collects large scale of single-image, multi-image, and video understanding data to train a more powerful vision language model. These two works appear after our work and incorporate our Mantis-Instruct as their training data, demonstrating our dataset’s value.
@@ 4.3 Multi-Image Evaluation
While many benchmarks have been developed for single-image tasks, such as TextVQA (Singh et al., 2019a), VQA-V2 (Goyal et al., 2016), MMBench (Liu et al., 2023d), and so on, the multi-image evaluation did not capture much attention until recently. Difference description is the first task which large-scale data and good evaluation benchmarks are developed for, including works like Spot-the-Diff (Jhamtani & Berg-Kirkpatrick, 2018), Birds-to-Words (Forbes et al., 2019), etc. They are usually evaluated by n-gram metrics like BLEU, and ROUGE, and only involve two images. The later NLVR2 (Suhr et al., 2018) further includes logical reasoning across 2 images into evaluation, focusing on a higher level of cognition ability. Q-bench (Wu et al., 2023a) and BLINK (Fu et al., 2024) are both proposed for low-level multi-image vision tasks, where humans easily do but model struggles. Another set of multi-image evaluations comes from video or image sequence evaluation, including Mementos (Wang et al., 2024b), MVBench (Li et al., 2023d), etc. In this case, the model usually needs to take more than 2 images to answer. Recently, there are also attempts to prompt LMMs to evaluate vision-language task (Zhang et al., 2023b), or image-generation tasks (Ku et al., 2023), which also demands good multi-image reasoning ability. The accuracies of the these tasks of different LMMs can be seen as another form of benchmark for multi-image ability.
@@ A.1 Details of Mantis-Instruct subsets
Co-Instruct (Wu et al., 2024): Co-Instruct is a dataset curated to compare the quality of 2-4 images in open-ended answer or multi-choice QA format. It’s similar to the tasks in the Qbench (Wu et al., 2023a).
@@
NLVR2 (Suhr et al., 2018): NLVR2 is a natural language reasoning dataset grounded in images. The task is to determine whether a statement is true or false given the visual contents of 2 images. NLVR2 covers many linguistic phenomena, such as cardinality, existential quantifiers, universal quantifiers, etc.
@@
Dreamsim (Fu et al., 2023): DreamSim proposes a dataset to train a model to judge image similarity. We reuse their datasets, and the task is to judge which candidate image is more similar to a reference image.
@@
ImageCoDe (Krojer et al., 2022): ImageCoDe is a multimodal challenge called image retrieval from contextual descriptions, where a multimodal model is required to retrieve/select the correct image from a set of 10 minimally contrastive images. For each image, 9 similar images are retrieved from Open Images Dataset V6 (Kuznetsova et al., 2018) using CLIP encodings.
@@
Contrast-Caption: Inspired by the paradigm of contrastive learning, we reformat existing image captioning tasks into a multi-image version, where 2 tasks are defined. Given multiple images, the first task we defined is to judge which image matches the provided caption, and the second task is to generate a caption for a denoted image. We randomly select 2 to 8 images along with their captions to form a single item, then apply the template of the 2 tasks defined. We used the images and captions provided by ShareGPT-4V (Chen et al., 2023) and LAION GPT-4V captions from LVIS (Schuhmann & Bevan, 2023).
@@
Spot-the-diff (Jhamtani & Berg-Kirkpatrick, 2018): This dataset requires the model to generate the difference description given 2 images. We use the processed version from MIMIC-IT (Li et al., 2023a).
@@
LRV-multi (Liu et al., 2023a): This is a single-image dataset used to mitigate hallucinations of LMMs. We process it similarly to LLaVA-665k-multi to get a "synthetic" multi-image dataset.
@@
Birds-to-Words (Forbes et al., 2019): The task of this dataset is to generate different descriptions and give 2 images of different birds. We directly take the training split for training. Besides, we prompt GPT-3.5-turbo to generate a possible question given the reference answer, so we can get a more diverse question pool instead of always asking the model to generate the difference description. Multi-VQA: This dataset contains multi-image reasoning QA pairs, where each question is generated by GPT-4 based on the captions of the images. Each question is required to involve at least two images. Images and captions are sampled from ShareGPT-4V (Chen et al., 2023). The prompt template is in  Table 2.
@@
Video QA: We also include some video question answering datasets, including NExTQA (Xiao et al., 2021), STAR (Wu & Yu, 2021), and visual-story-telling (Huang et al., 2016). nextqa and star are the processed version from Flipped-VQA (Ko et al., 2023). The visual-story-telling are the processed version from MIMIC-IT (Li et al., 2023a).
@@
Single-image-VQA: Instead of focusing only on multi-image tasks, we claim that it’s necessary to also include some single-image VQA dataset to avoid catastrophic forgetting on single-image tasks. We included some of the single image data from LLaVA-665k as well as DocVQA (Mathew et al., 2020), DVQA (Kafle et al., 2018), and ChartQA (Masry et al., 2022) to enhance its ability on diagrams and OCR. In practice, since the size of DVQA is too large, we only use 30k out of 200k examples for the training.
A.2 Investigation of Mantis’s capability on open-ended tasks
To assess Mantis’s performance on open-ended generation tasks, we evaluated it on image captioning (using the CoCo-2017-Lite dataset) (Chen et al., 2015) and video captioning (using the Vatex dataset) (Wang et al., 2019). For evaluation, we utilized lmms-eval Li* et al. (2024). The results are presented in  Table 9 and  Table 10. As shown in these tables, Mantis-Idefics2’s captioning performance improved after training on Mantis-Instruct compared to its previous checkpoint Idefics2, though still fall behind of the expert captioning models in these areas.
We attribute this improvement to the design of our dataset, which encourages the model to differentiate between images and better understand semantic information within visual content. Notably, contrastive learning of this kind is often limited in vision-language pre-training datasets, so fine-tuning on Mantis-Instruct appears to enhance the model’s comprehension of visual content further.
Model BLEU-4 BLEU-3 BLEU-2 BLEU-1 METEOR ROUGE-L CIDEr
Idefics2 0.1890 0.2810 0.4128 0.5896 0.2453 0.4545 0.7453
Mantis-Idefics2 0.2288 0.3266 0.4570 0.6155 0.2670 0.4906 0.8210
Mantis-SIGLIP-LLaMA3 0.2474 0.3479 0.4824 0.6500 0.2692 0.5045 0.9042
Table 9: Results of Mantis on the image captioning task (CoCo-2017-Lite)
Model BLEU-4 METEOR ROUGE-L CIDEr
Idefics2 0.1454 0.1448 0.3183 0.2307
Mantis-Idefics2 0.1702 0.1573 0.3724 0.2667
Mantis-SIGLIP-LLaMA3 0.1545 0.1431 0.3607 0.2263
Table 10: Results of Mantis on the video captioning task (Vatex)
A.3 Investigation of Mantis’s capability on long-context scenarios
While claiming Mantis to be a model that can accept text-image interleaved inputs, it’s interesting to investigate how Mantis will perform in long-context scenarios.
We first conduct experiments to see how Mantis’s performance will change as the number of images/frames increases. We run Mantis-Idefics2 on MVBench using 2/4/8/16 frames and report the performance in  Table 11. It turns out that Mantis‘s performance is stable across various numbers of sampled frames. What’s worth noting is that Mantis can use only 2 frames to achieve 48.38 accuracy, which means that most of the questions in the MVBench can be answered by keyframes. As the number of frames goes up from 2 to 8, the accuracy also goes up to 51.38 accuracy. The performance under 16 samples is 1 point lower than the performance under 8 frames, we hypothesize that the redundant information might distract Mantes from giving the current answer.
Number of frames 2 4 8 16
MVBench Acc 48.38 50.02 51.38 50.2
Table 11: Mantis-Idefics2’s performance on MVBench with different number of uniformly sampled frames.
As shown in  Table 4, MileBench consists of examples with an average of 15 images per example and a sequence length ranging from 1,000 to 9,000 tokens, making it well-suited for evaluating Mantis’s long-context ability. As shown in  Table 5, Mantis-SIGLIP achieves 47.5 on MileBench, only 5.5 points behind GPT-4V. We do believe this is good evidence that Mantis maintains competitive multimodal long-context ability.
A.4 Comparison with MIMIC-IT
MIMIC-IT is also a multi-modal dataset that aims to boost the model’s ability for in-context learning, which also contains multiple images for most of the QA pair (Li et al., 2023a). To investigate how well Mantis-Instruct outperforms MIMIC-IT, we conduct a fair comparison between the Otter and Mantis-Flamingo. As shown in  Table 12, Otter is tuned from the Open-Flamingo using MIMIC-IT as the instruction-tuning dataset, and one of our Mantis’s variants, Mantis-Flamingo, is also tuned from the Open-Flamingo but instead using our Mantis-Instruct as the instruction-tuning dataset. Therefore, Mantis-Flamingo and Otter are a fair comparison where they share the same pre-training dataset, MMC4+LAION, and trained on different instruction-following datasets, MIMIC-IT and Mantis-Instruct, respectively.
We report the performance in  Table 13. It turns out that Mantis-Flamingo is better than Otter in all these benchmarks, based on the same pre-trained model but different instruction-tuning datasets. This proves the superior quality of Mantis-Instruct compared to the MIMIC-IT. This is mainly because Mantis-Instruct is designed to equip models with the 4 essential abilities for multi-image reasoning, while MIMIC-IT is mainly designed for multimodal in-context reasoning.
Model Pre-training Instruction-Tuning Size
Otter MMC4, LAION MIMIC-IT 9B
Mantis-Flamingo MMC4, LAION Mantis-Instruct 9B
Table 12: Comparison of Mantis-Flamingo and Otter’s training datasets. Mantis-Flamingo and Otter are a fair comparison where they share the same pre-training dataset, MMC4+LAION, and trained on different instruction-following datasets, MIMIC-IT (Li et al., 2023a) and Mantis-Instruct, respectively
Model NVLR2 Q-Bench Mantis-Eval BLINK MVBench Average
Otter 49.15 17.50 13.29 36.26 15.30 26.30
Mantis-Flamingo 52.96 46.80 32.72 38.00 40.83 42.26
$\Delta$ over Otter +3.81 +29.30 +19.43 +1.74 +25.53 +15.96
Table 13: Comparison of Mantis-Flamingo and Otter’s (Li et al., 2023b) performance in multi-image benchmarks. The result numbers are sourced from  Table 5.